Pukri Ltd is deciding whether to pay out R90000jn excess cash in the form of an extra dividend or a share repurchase. Current profits are R2,40 per share and the share sells for R20. Balance sheet before paying out the dividend.
Equity 240 000. Dept 160 000. Bank/Cash 90 000. Other assets 310 000. Evaluate each alternative by
1.1 calculating the number of shares in issue
1.2 the dividends per share
1.3.1 calculate the new share price
1.3.2 calculate the EPS
1.3.3 The price earnings ratio
1
Expert's answer
2020-10-05T11:00:21-0400

Cash Dividend Option

NO of shares "= \frac{240000}{20} = 12000"

1.2 Dividend per share "= \frac{90000 }{12000} = 7.50"

1.3.1

New Share Price = 20 - 7.50 = 12.50

Note : - price per share decreases by the amount of dividend paid

1.3.2

EPS is unaffected because dividends are paid

EPS = 2.4

1.3.3

"PE Ratio = \frac{Price }{ Earnings}"

"PERatio = \frac{12.50 }{ 2.40} = 5.208"

Repurchase of Shares Option

NO of shares "= \frac{240000 }{ 20} = 12000"

shares repurchased "= \frac{90000 }{ 20} = 4500"

Remaining shares = 12000 - 4500 = 7500

1.2

There will be no dividends in repurchase option

1.3.1

Share Price will be same as before

Share Price = 20

1.3.2

Earnings "= 12000 \times 2.4 = 28800"

EPS = Earnings "= \frac{12.50 }{ 2.40} = 5.208"

New No of shares "= \frac{28800 }{ 7500} = 3.84"

1.3.3

PE Ratio "= \frac{Price }{ Earning} = \frac{20 }{ 3.84} = 5.208"